OpenStudy (anonymous):

the length of a rectangle is 3 inches more than its width. Write an equation for the situation above by multiplying the factors together and setting the equation equal to zero

OpenStudy (anonymous):

perhaps you know the area as well?

OpenStudy (anonymous):

sorry;{ 180=(w+3)(w)

OpenStudy (anonymous):

ah from before then \[w^2+3w=180\] \[w^2+3w-180=0\] \[(w+12)(w-15)=0\] so \[w=15\] because -12 makes no sense in this problem